OpenStudy (anonymous):

What is the possible discriminant of the graph? –13 0 15 16

OpenStudy (anonymous):



OpenStudy (anonymous):

(-7,0) and (-3,0)

OpenStudy (anonymous):

plz i need the answer my time is almost over and this is the only that i need

OpenStudy (asnaseer):

for a standard quadratic equation:\[ax^2+bx+c=0\]do you know what the discriminant is?

OpenStudy (anonymous):

I dont get it very well plz i just need the abswer this time im thinking but my time is almost over

OpenStudy (asnaseer):

the discriminant is defined as:\[b^2-4ac\]in your case you are given that the graph intersects the x-axis at (-7,0) and (-3,0). so you know the roots of the quadratic must be x=-7 and x=-3, which means it can be written as:\[(x+7)(x+3)=0\]

OpenStudy (asnaseer):

if you expand this out you get:\[x^2+10x+21=0\]so now you know: a = 1 b = 10 c = 21 plug those values into the formula for the discrimant

OpenStudy (anonymous):

thank you so much

OpenStudy (asnaseer):

yw

OpenStudy (anonymous):

is 16 right?

OpenStudy (asnaseer):

yes